OpenStudy (anonymous):

A Cold Stone Creamery ice cream shop sells sundaes for $3.60 and banana splits for $4.25. The shop sells four times as many sundaes as banana splits. If total sales amounted to $3,730 last weekend, how many banana splits were sold? What were the total sales of sundaes?

OpenStudy (anonymous):

Let banana spilit = x ,and undeas be 4x x*4.25 + 4x*3.60 = 3730 4.25x + 14.40x = 3730 18.65x = 3730 x = 3730/18.65 = 200 banana spilit = 200

OpenStudy (nurali):

banana spilit = x sundaes be 4x x(4.25) + 4x(3.60) = 3730 4.25x + 14.40x = 3730 18.65x = 3730 x = 3730/18.65 = 200 banana spilit = 200

OpenStudy (anonymous):

What were the total sales of sundaes?

OpenStudy (anonymous):

12